A drug company has a monopoly on a new patented medicine. The product can be made in either of two plants. The marginal costs of production for the two plants are MC1-30+20 MC2 10+302 P-30-3(+2) and The firms estimate of demand for the product is How much should the firm plan to produce in each plant? At what price should it plan to sell the product? (Round your responses to two decimal places.) The firm should produce units in plant 1 and n plant 2 To maximize profits, it should charge a price of $ per unit.
Answer #1

6 4 0 50505050 cold 6 7418 2221 Q012345 0369 012345 012345

From the above graph we can notice that the MC2 is above the demand curve which means that the cost of production is more. Hence we ignore MC2 and take MC1.

In order to determine profit maximizing level of output, we need to equate MC=MR. With an inverse linear demand curve, we know that the MR curve has the same vertical intercept but twice the slope, MR= 30-6Q

MC=MR
10+3Q=30-6Q
9Q=20
Q= 20/9= 2.22
P= 30-3*(2.22) =23.34

The firm should produce 0 units in plant 1 and 2.22 units in plant 2. To maximize profit it should charge $23.34 per unit